OpenStudy (anonymous):

HELP PLZZ!! becky leaves rapid city for laverne at 65mph. wendy leaves laverne for rapid city at 75mph. if rapid city and laverne are 420 miles apart, how long will it take becky and wendy to pass each other?

OpenStudy (biohazard9064):

It will take 3 hr

OpenStudy (anonymous):

U SURE?

OpenStudy (anonymous):

thank you <3

OpenStudy (biohazard9064):

yes

OpenStudy (anonymous):

420 / (75 + 65) = 420 / 140 = 3 hours

OpenStudy (anonymous):

thank you <3

sam (.sam.):

@BioHazard9064 Please explain the problem or guide the OP next time, rather than giving out answers.

OpenStudy (anonymous):

yea id prefer that coz i wanna learn but thanks for the help ether way.

OpenStudy (biohazard9064):

@.Sam. I said that but I was chastised for saying and doing that